# General approach for partitioning and phase separation in macromolecular coexisting phases

### v1

@@ I Introduction
cDFT relies on the minimization of a free energy functional. A central challenge lies in identifying or constructing a functional that accurately captures the physical behavior of the system under study. Over the past several decades, significant effort has been devoted to both the development of new functionals and the refinement of existing ones (for reviews see Evans et al. (2016); te Vrugt et al. (2020) and for recent developments, see e.g. Tschopp et al. (2020); Chen et al. (2024); Tschopp and Brader (2021); Sammüller et al. (2024, 2023)). Here, we combine fundamental measure theory (FMT) Rosenfeld (1989), free volume (FV) theory Oversteegen and Roth (2005) and mean field (MF) theory Archer et al. (2017) for obtaining the coexisting densities. In addition to using the standard MF (SMF) formalism as benchmark, we propose an enhanced MF (EMF) alternative that explicitly considers the effective free volume available to the different components. On the other hand, for obtaining inhomogeneous density profiles, we combine the Asakura-Osawa-Vrij (AOV) model Schmidt et al. (2000) given for the model polymer within the framework of FMT, with our EMF formalism. We demonstrate the applicability of our approach by tackling a representative case from aqueous two-phase systems used in water purification from pollutants Assis et al. (2021), as well as for biomaterial purification Iqbal et al. (2016), where two immiscible polymers phase-separate in the presence of a third dispersed material. More broadly, our findings can be extended to other systems, such as biomolecular condensates containing dispersed globular proteins or virus-like particles.
@@ II.1 Interaction potential
where $\sigma_{cp}=(\sigma_{pp}+\sigma_{cc})/2$. The second, describing surface attraction (e.g. due to electrostatic or solvophobic interactions), is defined as
@@ II.2 Free energy functional for bulk system
We start by defining the Helmholtz free energy functional per unit volume of a system with n polymers and m colloids Evans (1979). Generally, it can be written as the sum of the ideal and the excess contributions, i.e.
@@
For calculating the coexisting densities of a phase separating system, one can focus on expressing the free energy for homogeneous systems. In this case, $\mathcal{F}_{\rm HC}$ can be expressed using the Rosenfeld’s approximation Rosenfeld (1989), which is very accurate within low density regimes. The latter reads Roth (2010); Figueroa-Gerstenmaier et al. (2003)
@@
where $v_{cp}$ is the apparent volume arising from steric interactions between colloid $c$ and polymer $p$, given by $v_{cp}=\tfrac{\pi}{6}\sigma_{cp}^{3}$. Equation (17) is obtained by considering $\beta_{1}=\frac{1}{2}$ and $\beta_{2}=\frac{3}{8}$ in Eq. (16), while replacing $v_{cc}$ with $v_{cp}$. The latter assumption allows us to also take into account the size of the polymers explicitly. Under this assumption, the corresponding effective polymer density is
@@ II.3 Coexistence criteria
where $\mathcal{F}$ is the Helmholtz free energy per unit volume given in Eq. (6). On the other hand, the total pressure, in the absence of any external potential, can be written as Archer et al. (2024)
@@
This nested construction ensures that all species densities remain positive. The structure can be interpreted as a sequence of branching ratios, where each $x_{i}$ governs the split of remaining density among successive species. This representation is particularly well-suited for numerical optimization and phase coexistence solvers, as it transforms the constrained problem of non-negative, normalized densities into an unconstrained optimization over $\rho_{\rm t}$ and $\{x_{i}\}\in(0,1)$. The proposed reformulation helps in restricting the solver to find a disjoint solution using simple numerical solving modules, such as, e.g., hbr, freely available in python. The numerical approach consists in the following: one begins by generating a trial solution in terms of reduced density. This trial solution is then transformed into real densities, which are substituted into the coexistence equations. The values are iteratively updated until convergence is achieved toward a disjoint solution in the reduced variables. Once convergence is confirmed, the solution is converted back into real densities. A detailed description of these steps is provided in Section II of the supplementary information (SI).
@@
where free energy per particle, $f$, is expressed in terms of reduced densities ($\{\rho_{\rm t},x_{1},...,x_{\rm n+m-1}\}$ for an $\rm n+m$-component system). $\begin{subarray}{c}S\subseteq\{x_{i-1},\dots,x_{\rm n+m-1}\}\end{subarray}$ represents summation performed with all the combination for the number of species (except for zero species), along with all the permutation over species type, for the given number of species (here $\rm n+m$). However all the elements must belong to the set $\{x_{i-1},...,x_{\rm n+m-1}\}$. $w_{j}({x_{j}})$ is given by
@@ II.4 Grand potential for inhomogeneous system
$\Phi_{2}=\sum_{i=1,j=1}^{\rm n+m}\left(n_{1}^{i}n_{2}^{j}-\mathbf{n}_{V1}^{i}\cdot\mathbf{n}_{V2}^{j}\right)\frac{\partial^{2}\phi_{0}}{\partial n_{3}^{i}\partial n_{3}^{j}}$ (39)
@@
$\Phi_{3}=\frac{1}{8\pi}\sum_{\begin{subarray}{c}i=1,j=1\\ k=1\end{subarray}}^{\rm n+m}\Bigg(\frac{n_{2}^{i}n_{2}^{j}n_{2}^{k}}{3}-n_{2}^{i}\,\mathbf{n}_{V2}^{j}\cdot\mathbf{n}_{V2}^{k}+\frac{3}{2}\left(\mathbf{n}_{V2}^{i}\mathcal{T}^{j}\mathbf{n}_{V2}^{k}-\mathrm{tr}\left[\mathcal{T}^{i}\mathcal{T}^{j}\mathcal{T}^{k}\right]\right)\Bigg)\frac{\partial^{3}\phi_{0}}{\partial n_{3}^{i}\partial n_{3}^{j}\partial n_{3}^{k}},$ (40)
where $\phi_{0}$ represents the zero-dimensional free energy ($\beta\mathcal{F}_{0}$) given by $\beta\mathcal{F}_{0}=(1-\sum_{i=\rm n+1}^{\rm n+m}\eta_{i}-\sum_{j=1}^{\rm n}\eta_{j})\ln(1-\sum_{i=\rm n+1}^{\rm n+m}\eta_{i})-\sum_{i=\rm n+1}^{\rm n+m}\eta_{i}$. $n_{0}^{i},n_{1}^{i},n_{2}^{i},n_{3}^{i},\mathbf{n}_{V1}^{i},\mathbf{n}_{V2}^{i}$ and $\mathcal{T}$ are the weighted densities, which, for inhomogeneous cases, are given by
$n_{k}^{i}(\mathbf{r})=\int\omega_{k}^{i}(\mathbf{r^{\prime}-\mathbf{r}})\rho_{i}(\mathbf{r^{\prime}})d\mathbf{r^{\prime}},\quad k=0,\dots,3,V_{1},V_{2},$ (41)
@@
Here $R_{i}$ is the radius of the colloids, i.e. $R_{i}=\sigma_{ii}/2$, $\Theta(\mathbf{r}-R_{i})$ is the Heaviside step function, and $\delta(\mathbf{r}-R_{i})$ is the Dirac-delta function. The vector and tensor weight functions are defined as $\mathbf{\omega}_{V1}^{i}(\mathbf{r})=\omega_{1}^{i}(\mathbf{r})\mathbf{r}/r$, $\mathbf{\omega}_{V2}^{i}(\mathbf{r})=\omega_{2}^{i}(\mathbf{r})\mathbf{r}/r$ and $\omega_{\mathcal{T}}^{i}(\mathbf{r})=\omega_{2}^{i}(\mathbf{r})[\mathbf{r}\mathbf{r}/r^{2}-\mathcal{I}/3]$, where tr is the classical trace, $\rm{tr}(\omega_{\mathcal{T}}^{i})=0$ and $\mathcal{I}$ is the 3x3 identity matrix. The corresponding weighted densities are obtained in the same way as in Eq. (41), for both colloids and polymers. The details of the methodology and its application is given in the SI.
@@
Here, $i,j\in\{1,...,\rm n+m\}$ and $c\in\{\rm n+1,...,n+m\}$, and $\phi_{ij}$ represents both the polymer-polymer repulsive potential via Eq. (1) and the colloid-polymer attraction. For the latter, instead of using the square well form in Eq. (5), we use a parametrized Gaussian potential, as detailed in the SI, which remains continuous when solving the inhomogeneous densities in one dimension.
The inhomogeneous equilibrium density profile are obtained by minimizing the grand potential $\Omega$ with respect to the densities of the different species, providing the solution Roth (2010); Figueroa-Gerstenmaier et al. (2003)
@@ III.1 Bulk phase properties
Before investigating the effect on phase separation of different physical properties of this ternary system, we provide a comparison between cDFT and particle-based simulations (similar to those in Ref. 10). Specifically, for particle based simulation we set the polymer-polymer interaction to be gaussian potential, as in Eq. (1). For colloid-colloid and colloid-polymer b, we use a standard Lennard-Jones potential cut at its minimum, to simulate purely repulsive interactions. On the other hand, for the colloid-polymer a interaction, we consider two different Mie potentials with exponents u and v. Specifically, we use 482448-24, 241224-12 and 12612-6 potentials. For completeness, the $\rm u-v$ Mie potential is given as
$\phi_{ij}^{\rm sim}(r)=\dfrac{\rm u}{\rm u-v}\left(\dfrac{\rm u}{\rm v}\right)^{\rm v/(u-v)}\epsilon_{ij}\Bigg[\left(\frac{\sigma_{ij}}{r}\right)^{\rm u}-\left(\frac{\sigma_{ij}}{r}\right)^{\rm v}\Bigg].$ (46)
@@
Also in the corresponding cDFT system, polymer-polymer interactions are defined as in Eq. (1), and their parameters as for particle-based simulations (see top part of Table 1). When it comes to describing colloid–colloid and colloid–polymer interactions, a one-to-one benchmark is not straightforward. In fact, in particle-based simulations these interactions are soft, while cDFT considers hard-core repulsion plus eventually attraction. To solve this challenge, we set the size of the corresponding hard-core in cDFT following the Barker-Henderson Barker and Henderson (1967) (BH) perturbation theory. BH theory provides a mapping from a soft repulsive potential to a hard-core diameter, $\sigma_{\rm BH}$, which is given by
@@
To show the validity of the cDFT predictions on the phase response of the system, in Fig. 2(a) we vary the total polymer density, $\bar{\rho}_{\rm p}$, from 0.54 to 0.8 with $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}$, while keeping $\bar{\rho}_{\rm c}$ fixed at $0.09$. The predictions of our cDFT are bench-marked against particle-based simulations for all the three cases. Panel (a) shows a monotonic increase in partitioning as the total polymer density is increased. We also see that our improved MF approach performs better in all cases up to high colloidal density. Similarly, in panel (b) of Fig. 2, we vary the density of the colloids from $\bar{\rho}_{\rm c}=0.09$ to $0.4$ while keeping $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$ fixed. Interestingly, we see a decrease in partitioning after reaching a maximum value at around $\bar{\rho}_{\rm c}=0.2$. The presence of colloids increases the effective polymer density by reducing the available volume, which in turn lowers the free energy through enhanced polymer–colloid attraction. This effect, however, reaches a maximum, as the gain is compensated by increased colloid–colloid and polymer–polymer repulsive interactions, along with a further reduction in available volume that limits polymer–colloid contributions when colloids are too close. Also here, the cDFT predictions are in good agreement with simulations, up to high colloidal concentrations. We can see that, with the exception of case-I, the EMF approach performs better than its standard version. To investigate this specific case, the sensitivity of both MF approaches with respect to the value of $\sigma_{\rm BH}$ has been addressed (not shown here). We find that the EMF approach remains more anchored to the presented predictions, while the SMF approach deviates from the presented curve quite rapidly. All-in-all, we can conclude that the EMF free energy functional proposed here (Eq. (22)) performs better than its standard version.
@@
Figure 3: In (a), binodal curve for the ternary colloid-polymer mixture described by the following parameters: $R_{\rm aa}=R_{\rm bb}=R_{\rm ab}=\sqrt{2}\sigma$, $\epsilon_{\rm aa}=\epsilon_{\rm bb}=2.0,\epsilon_{\rm ab}=2.5$, $\sigma_{\rm aa}=\sigma_{\rm bb}=\sigma_{\rm cc}=1$ and $V_{\rm ac}=-5.6$. $x$ is the fraction of polymer b with reference to the total polymer densities $\bar{\rho}_{\rm p}$ and $y$ is the fraction of colloids with reference to the total density $\bar{\rho}_{\rm t}$. Here $\bar{\rho}_{\rm c}=0.09$ is kept fixed. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $\bar{\rho}_{\rm p}$. The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. The full circle represents the critical point. In (b), effect of polymer concentration on the coexistence densities of colloids.
@@
So far, we focused on the symmetric case $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}$. In order to investigate the effect of the relative amount of the two polymer species, we now fix the colloids density $\bar{\rho}_{\rm c}=0.09$, and one polymer species in turns. The results are shown in Fig. 5. Specifically, in panels (a) and (b), we fix $\bar{\rho}_{\rm b}=\rho_{\rm r}=0.27$ and vary $\bar{\rho}_{\rm a}$ around the reference value, $\rho_{\rm r}$. As one can see in panel (a), the difference between the coexisting densities of the colloids in the two phases remain almost unaltered (the tielines are practically parallel). This can also be seen in the inset of panel (b)). On the other hand, their value changes, and correlates with the value of $\bar{\rho}_{\rm a}$. The values of the coexisting densities of the polymers, however, depends on $\bar{\rho}_{\rm a}$, as expected. To compensate the change in the density of species a, a shift in $V_{\rm f}$ away from value of the symmetric case, $V_{f}=0.5$, is observed (see also inset of panel (a)). Figure 5(b) shows the variation in the number of particles, normalized by their total amount in the system, $N_{0}$, of the different system components as a function of $\bar{\rho}_{a}$ in the two phases. When $\bar{\rho}_{\rm a}$ is small, polymer a barely phase separate, yet partitioning is quickly increased as $\bar{\rho}_{a}$ is increased. Interestingly, there is a cross-over of the partitioning of the colloids. In fact, for small values of $\bar{\rho}_{a}$, colloids prefer phase-II. This trend is then reversed as $\bar{\rho}_{a}$ is increased. The partitioning of the colloids correlates with an increased volume of phase-I. Nevertheless, as already mention above, the coexisting densities of the colloids in the two phases vary only marginally (see inset of panel (b)). On the other hand, the relative amount of polymer b in the two phases is practically unaffected.

### v2

@@ I Introduction
cDFT relies on the minimization of a free energy functional. A central challenge lies in identifying or constructing a functional that accurately captures the physical behavior of the system under study. Over the past several decades, significant effort has been devoted to both the development of new functionals and the refinement of existing ones (for reviews see Evans et al. (2016); te Vrugt et al. (2020) and for recent developments, see e.g. Tschopp et al. (2020); Chen et al. (2024); Tschopp and Brader (2021); Sammüller et al. (2024, 2023)). Here, we combine fundamental measure theory (FMT) Rosenfeld (1989), free volume (FV) theory Oversteegen and Roth (2005) and mean field (MF) theory Archer et al. (2017) for obtaining the coexisting densities. In addition to using the standard MF (SMF) formalism as benchmark, we propose an enhanced MF (EMF) alternative that explicitly considers the effective free volume available to the different components. On the other hand, for obtaining inhomogeneous density profiles, we combine the Asakura-Osawa-Vrij (AOV) model Schmidt et al. (2000) given for the model polymer within the framework of FMT, with our EMF formalism. We demonstrate the applicability of our approach by tackling a representative case from aqueous two-phase systems used in water purification from pollutants Assis et al. (2021), as well as for biomaterial purification Iqbal et al. (2016), where two immiscible polymers phase-separate in the presence of a third dispersed material. Other relevant applications of these systems involve, e.g., the investigation of ultralow interfacial tension and micropattern formation in ferrofluids Rigoni et al. (2022), as well as their utilization in cell-mimicking microrobotic systems Furuki et al. (2024). More broadly, our findings can be extended to other systems, such as biomolecular condensates Holehouse and Alberti (2025) containing, e.g., dispersed globular proteins or virus-like particles.
@@ II.1 Interaction potential
where $\sigma_{cp}=(\sigma_{pp}+\sigma_{cc})/2$, where $\sigma_{pp}$ represents the excluded volume effect of the polymers in the presence of colloids (see discussion on free energy for hardcore interaction in Sec. II B). The second, describing surface attraction (e.g. due to electrostatic or solvophobic interactions), is defined as
@@ II.2 Free energy functional for bulk system
We start by defining the Helmholtz free energy functional per unit volume of a system with n polymers and m colloids. Generally, it can be written as the sum of the ideal and the excess contributions Evans (1979), i.e.
@@
For calculating the coexisting densities of a phase separating system, one can focus on expressing the free energy for homogeneous systems. In this case, $\mathcal{F}_{\rm HC}$ can be expressed using the Rosenfeld’s approximation Rosenfeld (1989), which is very accurate within low density regimes. The latter reads Roth (2010)
@@
where $v_{cp}$ is the apparent volume arising from steric interactions between colloid $c$ and polymer $p$, given by $v_{cp}=\tfrac{\pi}{6}\sigma_{cp}^{3}$. Equation (17) is obtained by considering $\beta_{1}=\frac{1}{2}$ and $\beta_{2}=\frac{3}{8}$ in Eq. (16), while replacing $v_{cc}$ (representing the volume of colloid $c$ in $\eta_{c}$) with $v_{cp}$. The latter assumption allows us to also take into account the size of the polymers explicitly. Under this assumption, the corresponding effective polymer density is
@@ II.3 Coexistence criteria
where $\mathcal{F}$ is the Helmholtz free energy per unit volume given in Eq. (6). On the other hand, the total pressure, in the absence of any external potential, can be written as
@@
This nested construction ensures that all species densities remain positive. The structure can be interpreted as a sequence of branching ratios, where each $x_{i}$ governs the split of remaining density among successive species. This representation is particularly well-suited for numerical optimization and phase coexistence solvers, as it transforms the constrained problem of non-negative, normalized densities into an unconstrained optimization over $\rho_{\rm t}$ and $\{x_{i}\}\in(0,1)$. The proposed reformulation helps in restricting the solver to find a disjoint solution using simple numerical solving modules, such as, e.g., hybr, freely available in python. The numerical approach consists in the following: one begins by generating a trial solution in terms of reduced density. This trial solution is then transformed into real densities, which are substituted into the coexistence equations. The values are iteratively updated until convergence is achieved toward a disjoint solution in the reduced variables. Once convergence is confirmed, the solution is converted back into real densities. A detailed description of these steps is provided in Section II of the supplementary information (SI).
@@
where free energy per particle, $f$, is expressed in terms of reduced densities ($\{\rho_{\rm t},x_{1},...,x_{\rm n+m-1}\}$ for an $\rm n+m$-component system). $\begin{subarray}{c}S\subseteq\{x_{i-1},\dots,x_{\rm n+m-1}\}\end{subarray}$ represents summation performed over all the possible subsets (all possible combination of different number of species and their permutation). However all the elements must belong to the set $\{x_{i-1},...,x_{\rm n+m-1}\}$. $w_{j}({x_{j}})$ is given by
@@ II.4 Grand potential for inhomogeneous system
$\Phi_{2}=\sum_{i=1,j=1}^{\rm n+m}\left(n_{1}^{i}n_{2}^{j}-\mathbf{n}_{\rm V1}^{i}\cdot\mathbf{n}_{\rm V2}^{j}\right)\frac{\partial^{2}\phi_{0}}{\partial n_{3}^{i}\partial n_{3}^{j}}$ (39)
@@
$\Phi_{3}=\frac{1}{8\pi}\sum_{\begin{subarray}{c}i=1,j=1\\ k=1\end{subarray}}^{\rm n+m}\Bigg(\frac{n_{2}^{i}n_{2}^{j}n_{2}^{k}}{3}-n_{2}^{i}\,\mathbf{n}_{\rm V2}^{j}\cdot\mathbf{n}_{\rm V2}^{k}+\frac{3}{2}\left(\mathbf{n}_{\rm V2}^{i}\mathcal{T}^{j}\mathbf{n}_{\rm V2}^{k}-\mathrm{tr}\left[\mathcal{T}^{i}\mathcal{T}^{j}\mathcal{T}^{k}\right]\right)\Bigg)\frac{\partial^{3}\phi_{0}}{\partial n_{3}^{i}\partial n_{3}^{j}\partial n_{3}^{k}},$ (40)
where $\phi_{0}$ represents the zero-dimensional free energy ($\beta\mathcal{F}_{0}$) given by $\beta\mathcal{F}_{0}=(1-\sum_{i=\rm n+1}^{\rm n+m}\eta_{i}-\sum_{j=1}^{\rm n}\eta_{j})\ln(1-\sum_{i=\rm n+1}^{\rm n+m}\eta_{i})-\sum_{i=\rm n+1}^{\rm n+m}\eta_{i}$. $n_{0}^{i},n_{1}^{i},n_{2}^{i},n_{3}^{i},\mathbf{n}_{\rm V1}^{i},\mathbf{n}_{\rm V2}^{i}$ and $\mathcal{T}$ are the weighted densities, which, for inhomogeneous cases, are given by
$n_{k}^{i}(\mathbf{r})=\int\omega_{k}^{i}(\mathbf{r^{\prime}-\mathbf{r}})\rho_{i}(\mathbf{r^{\prime}})d\mathbf{r^{\prime}},\quad k=0,\dots,3,\rm V1,V2,$ (41)
@@
Here $R_{i}$ is the radius of the colloids, i.e. $R_{i}=\sigma_{ii}/2$, $\Theta(\mathbf{r}-R_{i})$ is the Heaviside step function, and $\delta(\mathbf{r}-R_{i})$ is the Dirac-delta function. The vector and tensor weight functions are defined as $\mathbf{\omega}_{\rm V1}^{i}(\mathbf{r})=\omega_{1}^{i}(\mathbf{r})\mathbf{r}/r$, $\mathbf{\omega}_{\rm V2}^{i}(\mathbf{r})=\omega_{2}^{i}(\mathbf{r})\mathbf{r}/r$ and $\omega_{\mathcal{T}}^{i}(\mathbf{r})=\omega_{2}^{i}(\mathbf{r})[\mathbf{r}\mathbf{r}/r^{2}-\mathcal{I}/3]$, where tr is the classical trace, $\rm{tr}(\omega_{\mathcal{T}}^{i})=0$ and $\mathcal{I}$ is the 3x3 identity matrix. The corresponding weighted densities are obtained in the same way as in Eq. (41), for both colloids and polymers. The details of the methodology and its application is given in the SI.
@@
Here, $i,j\in\{1,...,\rm n+m\}$ and $c\in\{\rm n+1,...,n+m\}$, and $\phi_{ij}$ represents both the polymer-polymer repulsive potential via Eq. (1) and the colloid-polymer attraction. For the latter, instead of using the square well form in Eq. (5), we use a parametrized Gaussian potential, as detailed in the SI, which by being continuous provides a better choice for obtaining the numerical solutions of the inhomogeneous densities in one dimension.
The inhomogeneous equilibrium density profile are obtained by minimizing the grand potential $\Omega$ with respect to the densities of the different species, providing the solution Roth (2010)
@@ III.1 Bulk phase properties
Before investigating the effect on phase separation of different physical properties of this ternary system, we provide a comparison between cDFT and particle-based simulations (similar to those in Ref. 10). Specifically, we set the polymer-polymer interaction to be gaussian potential, as in Eq. (1). For colloid-colloid and colloid-polymer b, we use a standard Lennard-Jones potential cut at its minimum, to simulate purely repulsive interactions. On the other hand, for the colloid-polymer a interaction, we consider two different Mie potentials with exponents u and v. Specifically, we use 482448-24, 241224-12 and 12612-6 potentials. For completeness, the $\rm u-v$ Mie potential is given as
$\phi_{\rm ac}^{\rm sim}(r)=\dfrac{\rm u}{\rm u-v}\left(\dfrac{\rm u}{\rm v}\right)^{\rm v/(u-v)}\epsilon_{\rm ac}\Bigg[\left(\frac{\sigma_{\rm ac}}{r}\right)^{\rm u}-\left(\frac{\sigma_{\rm ac}}{r}\right)^{\rm v}\Bigg].$ (46)
@@
Also in the corresponding cDFT system, polymer-polymer interactions are defined as in Eq. (1), and their parameters as for particle-based simulations (see top part of Table 1). When it comes to describing colloid–colloid and colloid–polymer interactions, a one-to-one benchmark is not straightforward. In fact, in our particle-based simulations these interactions are soft, while cDFT considers hard-core repulsion plus eventually attraction. To solve this challenge, we set the size of the corresponding hard-core in cDFT following the Barker-Henderson Barker and Henderson (1967) (BH) perturbation theory. BH theory provides a mapping from a soft repulsive potential to a hard-core diameter, $\sigma_{\rm BH}$, which is given by
@@
To show the validity of the cDFT predictions on the phase response of the system, in Fig. 2(a) we vary the total polymer density, $\bar{\rho}_{\rm p}$, from 0.54 to 0.8 with $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}$, while keeping $\bar{\rho}_{\rm c}$ fixed at $0.09$. The predictions of our cDFT are bench-marked against particle-based simulations for all the three cases. Panel (a) shows a monotonic increase in partitioning as the total polymer density is increased. We also see that our improved MF approach performs better in all cases up to high colloidal density. Similarly, in panel (b) of Fig. 2, we vary the density of the colloids from $\bar{\rho}_{\rm c}=0.09$ to $0.4$ while keeping $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}=0.27$ fixed. Interestingly, we see a decrease in partitioning after reaching a maximum value at around $\bar{\rho}_{\rm c}=0.2$. The presence of colloids increases the effective polymer density by reducing the available volume, which in turn lowers the free energy through enhanced polymer–colloid attraction. This effect, however, reaches a maximum, as the gain is compensated by increased colloid–colloid and polymer–polymer repulsive interactions, along with a further reduction in available volume that limits polymer–colloid contributions when colloids are too close to each other. Also here, the cDFT predictions are in good agreement with simulations, up to high colloidal concentrations. We can see that, with the exception of case-I, the EMF approach performs better than its standard version. To investigate this specific case, the sensitivity of both MF approaches with respect to the value of $\sigma_{\rm BH}$ has been addressed. We find that the EMF approach remains more anchored to the presented predictions, while the SMF approach deviates from the presented curve quite rapidly. All-in-all, we can conclude that the EMF free energy functional proposed here (Eq. (22)) performs better than its standard version.
@@
Figure 3: In (a), binodal curve for the ternary colloid-polymer mixture described by the following parameters: $R_{\rm aa}=R_{\rm bb}=R_{\rm ab}=\sqrt{2}\sigma$, $\epsilon_{\rm aa}=\epsilon_{\rm bb}=2.0,\epsilon_{\rm ab}=2.5$, $\sigma_{\rm aa}=\sigma_{\rm bb}=\sigma_{\rm cc}=1$ and $\hat{V}_{\rm ac}=-5.6$. $x$ is the fraction of polymer b with reference to the total polymer densities $\bar{\rho}_{\rm p}$ and $y$ is the fraction of colloids with reference to the total density $\bar{\rho}_{\rm t}$. Here $\bar{\rho}_{\rm c}=0.09$ is kept fixed. The upper triangles indicate phase-I (polymer a-rich phase), while the lower triangles indicate phase-II (polymer b-rich phase). Coexisting points are connected by isobars, where their color corresponds to the value of $\bar{\rho}_{\rm p}$. The green squares represent the value of $V_{\rm f}$, i.e. the volume ratio of phase-I, while the red squares represent $V_{\rm f}=0.5$. The full circle represents the critical point. In (b), effect of polymer concentration on the coexistence densities of colloids.
@@
So far, we focused on the symmetric case $\bar{\rho}_{\rm a}=\bar{\rho}_{\rm b}$. In order to investigate the effect of the relative amount of the two polymer species, we now fix the colloids density $\bar{\rho}_{\rm c}=0.09$, and one polymer species in turns. The results are shown in Fig. 5, which illustrates a simple case of composition-driven phase separation, similar to those that can be observed in biological systems, such as, e.g., protein solutions (see, e.g., Farag et al. (2023)). Specifically, in panels (a) and (b), we fix $\bar{\rho}_{\rm b}=\rho_{\rm r}=0.27$ and vary $\bar{\rho}_{\rm a}$ around the reference value, $\rho_{\rm r}$. As one can see in panel (a), the difference between the coexisting densities of the colloids in the two phases remain almost unaltered (the tielines are practically parallel). This can also be seen in the inset of panel (b)). On the other hand, their value changes, and correlates with the value of $\bar{\rho}_{\rm a}$. The values of the coexisting densities of the polymers, however, depends on $\bar{\rho}_{\rm a}$, as expected. To compensate the change in the density of species a, a shift in $V_{\rm f}$ away from value of the symmetric case, $V_{f}=0.5$, is observed (see also inset of panel (a)). Figure 5(b) shows the variation in the number of particles, normalized by their total amount in the system, $N_{0}$, of the different system components as a function of $\bar{\rho}_{a}$ in the two phases. When $\bar{\rho}_{\rm a}$ is small, polymer a barely phase separate, yet partitioning is quickly increased as $\bar{\rho}_{a}$ is increased. Interestingly, there is a cross-over of the partitioning of the colloids. In fact, for small values of $\bar{\rho}_{a}$, colloids prefer phase-II. This trend is then reversed as $\bar{\rho}_{a}$ is increased. The partitioning of the colloids correlates with an increased volume of phase-I. Nevertheless, as already mention above, the coexisting densities of the colloids in the two phases vary only marginally (see inset of panel (b)). On the other hand, the relative amount of polymer b in the two phases is practically unaffected.